OpenStudy (anonymous):

A parallelogram has an area of 396 square centimeters and a height of 18 centimeters. What is the length of the base of the parallelogram? cm

OpenStudy (anonymous):

396 square cms divided by 18 cms is the answer.

OpenStudy (texaschic101):

A = B * H A = 396 H = 18 now we sub 396 = 18B --- divide both sides by 18 396/18 = B you can do the math now...

OpenStudy (anonymous):

22

OpenStudy (anonymous):

Correct!